OpenStudy (anonymous):

what is the radius of the circle shown by: x^2+y^2-2x+6y+9=0

OpenStudy (ranga):

Group the x terms together and complete the square. Group the y terms together and complete the square. Put it in the form: (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center of the circle and r is the radius.

OpenStudy (anonymous):

I got a radius of 1, is that correct? @ranga

OpenStudy (ranga):

x^2+y^2-2x+6y+9=0 (x^2 - 2x) + (y^2 + 6y) + 9 = 0 (x-1)^2 - 1 + (y+3)^2 - 9 + 9 = 0 (x-1)^2 + (y+3)^2 = 1 Yes, the radius is 1. Good job!